OpenStudy (anonymous):

Simplify completely quantity x squared plus x minus 12 over quantity x squared minus x minus 20

OpenStudy (anonymous):

here is the question in math terms

OpenStudy (anonymous):

can someone please help me with this i can't figure it out.

OpenStudy (anonymous):

anyone? please help

OpenStudy (anonymous):

x^2 +x-12\[\frac{ x^2 +x-12 }{ x^2-x-20 }\]

OpenStudy (anonymous):

factor the numerator and denominator

OpenStudy (anonymous):

\[\frac{ (x-3)(x+4) }{ (x-5)(x+4) }\]

OpenStudy (anonymous):

x+4 cancels out

OpenStudy (anonymous):

left with (x-3)/(x-5)

OpenStudy (anonymous):

these are my only choices

OpenStudy (anonymous):

\[3\frac{ x^2-8x+15 }{ 4x^2-16x-20 }\]

OpenStudy (anonymous):

this factors to \[3\frac{ (x-5)(x-3) }{4(x-1)(x-5) }\]

OpenStudy (anonymous):

cancel out the x-5s

OpenStudy (anonymous):

left with \[\frac{ 3(x-3) }{ 4(x-1)}\]

OpenStudy (anonymous):

now divide this term by multiplying the reciprocal

OpenStudy (anonymous):

\[\frac{ x-3 }{ x-5 }*\frac{ 4(x-1) }{ 3(x-3) }\]

OpenStudy (anonymous):

x-3s cancel out leaving you with:\[\frac{ 4(x-1) }{ x-5 }\] that's completely simplified

OpenStudy (anonymous):

the only choice i have like why you put is 4(x+1) not what u put

OpenStudy (anonymous):

dunno then